# A Local Projection Stabilised HHO Method for the Oseen Problem

### v1

@@ 4 A Priori Error Estimates
This section deals with the a priori error analysis for the discrete solution of velocity and pressure from (2.7). We employ the approximation results in (1.4) to compute the a priori error under the $\left|\!\left|\!\left|\cdot\right|\!\right|\!\right|_{{\rm LP}}$ norm.
@@
Let $(\boldsymbol{u},p)\in\boldsymbol{V}\times Q$ be the solution to the continuous problem (1.2) and $(\underline{\boldsymbol{u}}_{h},p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$ be the solution to the discrete problem (2.7). Assume that $\boldsymbol{u}\in\boldsymbol{V}\cap[H^{k+2}(\mathcal{T}_{h})]^{2}$ such that $\Delta\boldsymbol{u}\in L^{2}(\Omega)$, and $\boldsymbol{u}\in H^{k+1}(\mathcal{M}_{h})$. Assume $p\in H^{1}(\Omega)\cap H^{k+1}(\mathcal{M}_{h})$. Then, the following result holds
@@
For simplicity of notation, set $(\tilde{\underline{\boldsymbol{u}}}_{h},\tilde{p}_{h})=(\underline{\boldsymbol{I}}_{h}^{k}\boldsymbol{u},\pi_{h}^{k}p)\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$. Then the error $(\tilde{\underline{\boldsymbol{u}}}_{h}-\underline{\boldsymbol{u}}_{h},\tilde{p}_{h}-p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$. Now applying Theorem 3.4 we have
@@
We estimate each of the terms in the above bilinear form $A_{h}^{LP}$. Using the definition of discrete problem (2.7), we have
@@
Since $\Delta\boldsymbol{u}\in[L^{2}(\Omega)]^{2}$ and $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{h,0}^{k}$, the following identity holds
@@
Now, using the expression of $(\boldsymbol{f},\boldsymbol{v}_{h})$ in (4.4) and the definition of $A_{h}^{LP}((\tilde{\underline{\boldsymbol{u}}}_{h},\tilde{p}_{h}),(\underline{\boldsymbol{v}}_{h},q_{h}))$, (4.3) can be rewritten as
@@
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon(\nabla\boldsymbol{u},\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}\epsilon(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{u}\cdot\boldsymbol{n})_{F}-A_{\epsilon,T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}$
@@
We estimate each of the above five terms of (4) starting with the diffusion consistency term $E_{1}$. Using the definition of the reconstruction operator $\boldsymbol{r}_{T}^{k+1}$ in (2.1) with $\boldsymbol{w}=\nabla r_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T}$ we have
$(\nabla\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{v}}_{T},\nabla\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})_{T}=(\nabla\boldsymbol{v}_{T},\nabla\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T}\cdot\boldsymbol{n}_{TF})_{F}.$ (4.6)
@@
$\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon(\nabla\boldsymbol{u},\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}\epsilon(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{u}\cdot\boldsymbol{n})_{F}-A_{\epsilon,T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}$
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T}),\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})\cdot\boldsymbol{n},\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}-S_{\epsilon,T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}.$ (4.7)
The velocity reconstruction operator satisfies $\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{I}}_{T}^{k}\boldsymbol{u}=\pi_{T}^{1,k+1}\boldsymbol{u}$, see (Piet_Jero_HHO_Book_20, , defn 1.39). Using this, the first term inside the summation of (4) vanishes. The second term of (4) can be controlled using the approximation property of $\boldsymbol{r}_{T}^{k+1}$ in (2.2) along with the Cauchy–Schwarz inequality and equivalence of the norms $\left\|\cdot\right\|_{\epsilon,h}$ and $\left\|\cdot\right\|_{\rm dG}$ as
$\sum_{F\in\mathcal{F}_{T}}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})\cdot\boldsymbol{n},\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}\leq C\epsilon^{1/2}h_{T}^{(k+1)}\lvert\boldsymbol{u}\rvert_{k+2,T}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\epsilon,T}.$
@@
$\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})\cdot\boldsymbol{n},\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}\epsilon h_{T}^{2(k+1)}\lvert\boldsymbol{u}\rvert^{2}_{k+2,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\epsilon,h}.$ (4.8)
Since $S_{\epsilon,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\tilde{\underline{\boldsymbol{u}}}_{h})\leq C\sum_{T\in\mathcal{T}_{h}}\epsilon h_{T}^{2(k+1)}\lvert\boldsymbol{u}\rvert^{2}_{k+2,T}$ (see Oseen_HHO ), the third term of (4) can be controlled by using the definition of $S_{\epsilon,h}$ and the Cauchy-Schwarz inequality
@@
Now we estimate the term $E_{2}$ in (4). Applying an integration by parts on $(\boldsymbol{b}\cdot\nabla\boldsymbol{u},\boldsymbol{v}_{T})_{T}$, using the definition of $A_{\boldsymbol{b},T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})$ and the fact that $((\boldsymbol{b}\cdot\boldsymbol{n})\left[\hskip-3.5pt\left[\boldsymbol{u}\right]\hskip-3.5pt\right],\boldsymbol{v}_{F})_{F}=0$ for all $F\in\mathcal{F}_{h}$ we have
@@
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}-(\boldsymbol{u},\boldsymbol{b}\cdot\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n})\boldsymbol{u},\boldsymbol{v}_{T})_{F}+\sigma(\boldsymbol{u},\boldsymbol{v}_{T})_{T}-A_{\boldsymbol{b},T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}$
$=\sum_{T\in\mathcal{T}_{h}}-(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T},\boldsymbol{b}\cdot\nabla\boldsymbol{v}_{T})_{T}+\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}\cdot\boldsymbol{n}(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T}),\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}$
$\quad-\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\Big{(}(\boldsymbol{b}\cdot\boldsymbol{n})^{\ominus}(\tilde{\boldsymbol{u}}_{F}-\tilde{\boldsymbol{u}}_{T}),(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\Big{)}_{F}+\sum_{T\in\mathcal{T}_{h}}\sigma(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T},\boldsymbol{v}_{T})_{T}.$ (4.11)
Let $\boldsymbol{b}_{T}$ be a $P_{0}$ approximation of $\boldsymbol{b}$ on $T$. Since $\pi_{T}^{k}$ is the $L^{2}$ projection, $(\boldsymbol{b}_{T}\cdot\nabla\boldsymbol{v}_{T},\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T})_{T}=0$ for all $T\in\mathcal{T}_{h}$. We subtract this from the first term in (4) and use (1.4) to get
@@
$\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+2}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b},h}.$ (4.12)
Using the approximation estimates in (1.4), the Cauchy–Schwarz inequality and the trace inequality, the second term in (4) gives
$\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}\cdot\boldsymbol{n}(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T}),\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}$ $\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\Big{(}\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{\lvert\boldsymbol{b}\cdot\boldsymbol{n}\rvert}{2}\left\|\boldsymbol{v}_{F}-\boldsymbol{v}_{T}\right\|_{F}^{2}\Big{)}^{1/2}$
$\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b},h}.$ (4.13)
The third term of (4) is the upwind stabilisation term which can be controlled in a manner similar to (4) as
$\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\Big{(}(\boldsymbol{b}\cdot\boldsymbol{n})^{\ominus}(\tilde{\boldsymbol{u}}_{F}-\tilde{\boldsymbol{u}}_{T}),(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\Big{)}_{F}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b},h}.$ (4.14)
@@
$\sigma(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T},\boldsymbol{v}_{T})_{T}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2(k+1)}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b},h}.$ (4.15)
Combining the estimates (4)–(4.15)we get
$E_{2}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b},h}.$ (4.16)
The third term of the consistency error in (4) is $E_{3}=(\boldsymbol{v}_{h},\nabla p)-B_{h}(\underline{\boldsymbol{v}}_{h},\tilde{p}_{h})$. Using the definition of the bilinear form $B_{h}$ and using $(\nabla\cdot\boldsymbol{v}_{T},\pi_{T}^{k}p)_{T}=(\nabla\cdot\boldsymbol{v}_{T},p)_{T}$ we arrive at
@@
$A_{S,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\underline{\boldsymbol{v}}_{h})=\tau_{M}(G_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{I}}_{T}^{k}(\boldsymbol{u})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{v}}_{h})))_{M}.$

### v2

@@ 4 A Priori Error Estimates
This section deals with the a priori error analysis for the discrete solution of velocity and pressure from (2.7). We employ the approximation results in (1.4) to compute the a priori error under the $\left|\!\left|\!\left|\cdot\right|\!\right|\!\right|_{{\rm LP}}$ norm. We follow various steps in consistency and convergence that are based on the 3rd Strang Lemma; see [25, 1].
@@
Let $(\boldsymbol{u},p)\in\boldsymbol{V}\times Q$ be the solution to the continuous problem (1.2) and $(\underline{\boldsymbol{u}}_{h},p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$ be the solution to the discrete problem (2.7). Assume that $\boldsymbol{u}\in\boldsymbol{V}\cap[H^{k+2}(\mathcal{T}_{h})]^{d}$ such that $\Delta\boldsymbol{u}\in[L^{2}(\Omega)]^{d}$, and $\boldsymbol{u}\in H^{k+1}(\mathcal{M}_{h})$. Assume $p\in H^{1}(\Omega)\cap H^{k+1}(\mathcal{M}_{h})$. Then, the following result holds
@@
For simplicity of notation, set $(\tilde{\underline{\boldsymbol{u}}}_{h},\tilde{p}_{h}):=(\underline{\boldsymbol{I}}_{h}^{k}\boldsymbol{u},\pi_{h}^{k}p)\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$. Then the error is $(\tilde{\underline{\boldsymbol{u}}}_{h}-\underline{\boldsymbol{u}}_{h},\tilde{p}_{h}-p_{h})\in\underline{\boldsymbol{U}}_{h,0}^{k}\times P_{h}^{k}$. Now applying Theorem 3.4, we have
@@
We estimate each of the terms in the above bilinear form $A_{h}^{LP}$. Using the definition of discrete problem (2.7), we have the following consistency error:
@@
Since $\Delta\boldsymbol{u}\in[L^{2}(\Omega)]^{d}$ and $\underline{\boldsymbol{v}}_{h}\in\underline{\boldsymbol{U}}_{h,0}^{k}$, the following identity holds
@@
Now, using the expression $(\boldsymbol{f},\boldsymbol{v}_{h})$ in (4.4) and the definition of $A_{h}^{LP}((\tilde{\underline{\boldsymbol{u}}}_{h},\tilde{p}_{h}),(\underline{\boldsymbol{v}}_{h},q_{h}))$, (4.3) can be rewritten as
@@
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon(\nabla\boldsymbol{u},\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}\epsilon(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{u}\cdot\boldsymbol{n}_{TF})_{F}-A_{\epsilon,T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}$
@@
We estimate each of the five above terms of (4) starting with the diffusion consistency term $E_{1}$. Using the definition of the reconstruction operator $\boldsymbol{r}_{T}^{k+1}$ in (2.1) with $\boldsymbol{w}=\nabla r_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T}$, we have
$(\nabla\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{v}}_{T},\nabla\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})_{T}=(\nabla\boldsymbol{v}_{T},\nabla\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T}\cdot\boldsymbol{n}_{TF})_{F}.$ (4.6)
@@
$\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon(\nabla\boldsymbol{u},\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}\epsilon(\boldsymbol{v}_{F}-\boldsymbol{v}_{T},\nabla\boldsymbol{u}\cdot\boldsymbol{n}_{TF})_{F}-A_{\epsilon,T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}$
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T}),\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})\cdot\boldsymbol{n}_{TF},\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}-S_{\epsilon,T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}.$ (4.7)
The velocity reconstruction operator satisfies $\boldsymbol{r}_{T}^{k+1}\underline{\boldsymbol{I}}_{T}^{k}\boldsymbol{u}=\pi_{T}^{1,k+1}\boldsymbol{u}$, see [25, Definition 1.39]. Using this, the first term within the summation of (4) vanishes. The second term of (4) can be controlled using the approximation property of $\boldsymbol{r}_{T}^{k+1}$ in (2.2) along with the Cauchy–Schwarz inequality and equivalence of the norms $\left\|\cdot\right\|_{\epsilon,h}$ and $\left\|\cdot\right\|_{1,h}$ as
$\sum_{F\in\mathcal{F}_{T}}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})\cdot\boldsymbol{n}_{TF},\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}\leq C\epsilon^{1/2}h_{T}^{(k+1)}\lvert\boldsymbol{u}\rvert_{k+2,T}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\epsilon,T}.$
@@
$\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\epsilon(\nabla(\boldsymbol{u}-\boldsymbol{r}_{T}^{k+1}\tilde{\underline{\boldsymbol{u}}}_{T})\cdot\boldsymbol{n}_{TF},\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}\epsilon h_{T}^{2(k+1)}\lvert\boldsymbol{u}\rvert^{2}_{k+2,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\epsilon,h}.$ (4.8)
Since $S_{\epsilon,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\tilde{\underline{\boldsymbol{u}}}_{h})\leq C\sum_{T\in\mathcal{T}_{h}}\epsilon h_{T}^{2(k+1)}\lvert\boldsymbol{u}\rvert^{2}_{k+2,T}$ (see [1]), the third term of (4) can be controlled by using the definition of $S_{\epsilon,h}$ and the Cauchy-Schwarz inequality as
@@
Now we estimate the term $E_{2}$ in (4). Applying an integration by parts on $(\boldsymbol{b}\cdot\nabla\boldsymbol{u},\boldsymbol{v}_{T})_{T}$, using the definition of $A_{\boldsymbol{b},T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})$ and the fact that $\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n}_{TF})\boldsymbol{u},\boldsymbol{v}_{F})_{F}=0$, we have
@@
$=\sum_{T\in\mathcal{T}_{h}}\Big{(}-(\boldsymbol{u},\boldsymbol{b}\cdot\nabla\boldsymbol{v}_{T})_{T}+\sum_{F\in\mathcal{F}_{T}}((\boldsymbol{b}\cdot\boldsymbol{n}_{TF})\boldsymbol{u},\boldsymbol{v}_{T})_{F}+\sigma(\boldsymbol{u},\boldsymbol{v}_{T})_{T}-A_{\boldsymbol{b},T}(\tilde{\underline{\boldsymbol{u}}}_{T},\underline{\boldsymbol{v}}_{T})\Big{)}$
$=\sum_{T\in\mathcal{T}_{h}}-(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T},\boldsymbol{b}\cdot\nabla\boldsymbol{v}_{T})_{T}+\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}\cdot\boldsymbol{n}_{TF}(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T}),\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}$
$\quad-\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\Big{(}(\boldsymbol{b}\cdot\boldsymbol{n}_{TF})^{\ominus}(\tilde{\boldsymbol{u}}_{F}-\tilde{\boldsymbol{u}}_{T}),(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\Big{)}_{F}+\sum_{T\in\mathcal{T}_{h}}\sigma(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T},\boldsymbol{v}_{T})_{T}.$ (4.11)
Let $\boldsymbol{b}_{T}$ be a $P_{0}$ approximation of $\boldsymbol{b}$ on $T$. Since $\pi_{T}^{k}$ is the $L^{2}$ projection, $(\boldsymbol{b}_{T}\cdot\nabla\boldsymbol{v}_{T},\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T})_{T}=0$ for all $T\in\mathcal{T}_{h}$. We subtract this from the first term in (4) and use (1.4), inverse inequality and $\boldsymbol{b}-\boldsymbol{b}_{T}\sim h_{T}$ to get
@@
$\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+2}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b}}.$ (4.12)
Using the approximation estimates in (1.4), the Cauchy–Schwarz inequality, and the trace inequality, the second term in (4) gives
$\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}(\boldsymbol{b}\cdot\boldsymbol{n}_{TF}(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T}),\boldsymbol{v}_{F}-\boldsymbol{v}_{T})_{F}$ $\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\Big{(}\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\frac{1}{2}\left\|\boldsymbol{b}\cdot\boldsymbol{n}_{TF}(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\right\|_{F}^{2}\Big{)}^{1/2}$
$\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b}}.$ (4.13)
The third term of (4) is the upwind stabilisation term which can be controlled using the estimate
$\|\tilde{\boldsymbol{u}}_{F}-\tilde{\boldsymbol{u}}_{T}\|_{F}=\|\pi_{F}^{k}\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}\|_{F}=\|\pi_{F}^{k}(\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u})\|_{F}\leq\|\boldsymbol{u}-\pi_{T}^{k}\boldsymbol{u}\|_{F}$
and steps similar to (4) as
$\sum_{T\in\mathcal{T}_{h}}\sum_{F\in\mathcal{F}_{T}}\Big{(}(\boldsymbol{b}\cdot\boldsymbol{n}_{TF})^{\ominus}(\tilde{\boldsymbol{u}}_{F}-\tilde{\boldsymbol{u}}_{T}),(\boldsymbol{v}_{F}-\boldsymbol{v}_{T})\Big{)}_{F}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b}}.$ (4.14)
@@
$\sigma(\boldsymbol{u}-\tilde{\boldsymbol{u}}_{T},\boldsymbol{v}_{T})_{T}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2(k+1)}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b}}.$ (4.15)
Combining the estimates (4)–(4.15), we get
$E_{2}\leq C\Big{(}\sum_{T\in\mathcal{T}_{h}}h_{T}^{2k+1}\lvert\boldsymbol{u}\rvert^{2}_{k+1,T}\Big{)}^{1/2}\left\|\underline{\boldsymbol{v}}_{h}\right\|_{\boldsymbol{b}}.$ (4.16)
The third term of the consistency error in (4) is $E_{3}=(\boldsymbol{v}_{h},\nabla p)-B_{h}(\underline{\boldsymbol{v}}_{h},\tilde{p}_{h})$. Using the definition of the bilinear form $B_{h}$ and using $(\nabla\cdot\boldsymbol{v}_{T},\pi_{T}^{k}p)_{T}=(\nabla\cdot\boldsymbol{v}_{T},p)_{T}$, we arrive at
@@
$A_{S,h}(\tilde{\underline{\boldsymbol{u}}}_{h},\underline{\boldsymbol{v}}_{h})=\tau_{M}(G_{\boldsymbol{b}_{M},M}(\underline{\boldsymbol{I}}_{T}^{k}(\boldsymbol{u})),K_{M}(\boldsymbol{G}_{\boldsymbol{b}_{M},M}^{k}(\underline{\boldsymbol{v}}_{h})))_{M}.$